OpenStudy (anonymous):

This may be a long question, but please bear with me. "use the vertex formula to find the vertex of the graph of f(x)=1.5-4x-12x^2." I got as far as finding x=-1/6. When I plugged that back into the original equation, I got completely lost as far as what to do with 1.5+2/3-1/3. This is the result after squaring. The textbook just says simplify but when I did that I got the wrong answer. Please help.

OpenStudy (anonymous):

Which grade are you in? this is a calculus based method\[\frac{dy}{dx}=\frac{d(1.5-4x-12x^2)}{dx}= -4-24x\]putting dy/dx=0, we get \[x = \frac{-1}{6}\]this is the vertex and can be checked by putting into the expression for y.\[y=1.5+\frac{4}{6}-12(\frac{-1}{6})^2= 1.5+2/3-1/3= 1.5+1/3 = 11/6\]

OpenStudy (anonymous):

Ah, I see my mistake. Thanks!